COUPLED CLUSTER CALCULATIONS FOR COMPLEXES BETWEEN AN ARGON ATOM AND A LINEAR MOLECULE OR CATION

\underline{PETER BOTSCHWINA} and RAINER OSWALD, Institut für Physikalische Chemie, Universität Göttingen, Tammanstraße 6, D-37077 Göttingen, Germany.

Geometric structures and various spectroscopic properties were calculated by the coupled cluster variant CCSD(T) for a number of complexes between an argon atom and linear neutral molecules (HCCH, HC_{3}N and HC_{4}H) and cations (HCNH^{+}, HCCH^{+}, HC_{3}NH^{+} and HC_{4}H^{+}). The electronic ground state of Ar⋯HCCH^{+} has ^{2}A_{1} symmetry with a T shaped equilibrium structure. The linear form of Ar···HC_{4}H^{+} is energetically favoured over the T form, while the opposite holds for the corresponding neutral complex.